# A Study on Chinese Social Perspective regarding ChatGPT for Education and Beyond

### v3

@@ Abstract
ChatGPT has piqued the interest of many fields, particularly in the academic community. GPT-4, the latest version, starts supporting multimodal input and output. This study examines social media posts to analyze how the Chinese public perceives the potential of ChatGPT for educational and general purposes. The study also serves as the first effort to investigate the changes in public opinion since the release of GPT-4. According to the analysis results, prior to GPT-4, although some social media users believed that AI advancements would benefit education and society, some believed that advanced AI, such as ChatGPT, would make humans feel inferior and lead to problems such as cheating and a decline in moral principles, while the majority remain neutral. Interestingly, public attitudes have tended to shift in a positive direction since the release of GPT-4. We present a thorough analysis of the trending shift and a roadmap to ensure the ethical application of ChatGPT-like models in education and beyond.
@@
Artificial Intelligence (AI), ChatGPT, Large Language Model (LLM), Chatbots, Education, Ethics, Human-Computer Interaction.I Introduction
Chatbots can understand natural language input and respond in a human-like manner, making them ideal for tasks such as answering questions, providing guidance, and even offering emotional support. In customer service, chatbots can handle common problems and provide support and guidance. For example, Cui et al. [1] introduced SuperAgent, which is cost-effective when answering repetitive questions, freeing up human support staff to answer more complicated questions. Chatbot can also help users find interesting products and recommend related products in online shopping services. As confirmed by [2], anthropomorphic bot plays a positive role in shaping consumers’ intentions to purchase through chatbot commerce. In the medical field, chatbots serve as medical manuals to help patients become aware of their illnesses and improve their health. For example, text diagnosis bot enables sufferers to join in analyses of their medicinal matters and present a personalized analysis report about the symptoms [3]. Meanwhile, research on the application of chatbots to education and other areas is still in its infancy [4] and waits for more exploration.
@@
We gathered data by crawling posts containing “ChatGPT” and other relevant keywords on Sina Weibo, China’s biggest microblogging website and app with 605 million monthly active users. Subsequently, we categorized the posts gathered by the specified keywords into two major groups: AI for education and human-AI interaction. Each major category includes three scenarios. We conducted sentiment and topic analyses on these posts to assess the positive and negative impacts of ChatGPT, aiming to enhance its benefits while mitigating potential drawbacks for its safe and responsible use.
@@
Public concerns include privacy infringement, content bias, inaccuracy, harm to educational fairness, and over-reliance on AI application.
@@ II Related Work
Papers | Brief Introduction |
Mogavi et al. [10] |
This article analyzed data from X, Reddit, YouTube, and LinkedIn to explore the user experience and perspectives of early adopters toward ChatGPT in various education sectors. |
Baidoo-Anu et al. [11] |
This review synthesized recent extant literature to offer some potential benefits of ChatGPT in promoting teaching and learning. |
Kasneci et al. [12] |
This commentary discusses the potential benefits and challenges of LLMs in education from student and teacher perspectives. |
Rudolph et al. [13] |
This review adopts a desktop analysis approach and conducts an extensive literature review, focusing on ChatGPT’s implications for higher education. |
Zhai [14] |
This study piloted ChatGPT to write an academic paper and reflected on the potential impacts of AI tools on education based on the user experience. |
Tlili et al. [15] |
This study collected and analyzed X data, and examined ChatGPT in education among early adopters through a qualitative instrumental case study. |
Kung et al. [16] |
This article evaluated the performance of ChatGPT on the United States Medical Licensing Exam and suggested that LLMs may have the potential to assist with medical education and clinical decision-making. |
Rahman et al. [17] |
This study analyzed data from articles, websites, blogs, and various media, finding ChatGPT effective in generating initial academic research ideas. However, it identified challenges in literature synthesis, citations, problem statements, research gaps, and data analysis, suggesting the need for guidelines on the proper use of LLMs. |
Bahrini et al. [18] |
This article reviewed the existing literature, and examined the applications, opportunities, and threats of ChatGPT in 10 main domains, providing detailed examples for the business and industry as well as education. |
Liu et al. [19] |
This article performed an in-depth analysis of 194 relevant papers on arXiv, and the findings reveal a significant and increasing interest in ChatGPT/GPT-4 research. |
Our article |
Our article differs in the following ways: 1) we collect data from Chinese social media, Weibo, which serves as a unique public opinion environment and data source of Chinese society that has not been explored; 2) we compare the people’s attitudes and opinions before and after the release of GPT-4; 3) we expand the investigation to more interactions scenarios beyond education. |
@@
Education is one of the most widely engaging application fields for ChatGPT, with a broad audience and easy deployment. Mogavi et al. [10] analyzed social media data from X, Reddit, YouTube, and LinkedIn, revealing that the public discourse is generally positive, and there is enthusiasm regarding ChatGPT’s use in education. Nevertheless, they found that an overreliance on AI systems may encourage superficial learning habits and diminish students’ social and critical thinking skills. Baidoo-Anu et al. [11] highlight ChatGPT’s capability to assess students’ learning styles, offer personalized tutoring and feedback, and even grade assignments, thereby freeing up teachers for other tasks. However, concerns about the absence of human interaction in such generative models are noted, especially for students needing personal teacher engagement for an optimal learning experience. Kasneci et al. [12] believe that ChatGPT can provide customized services for learners at different stages and fields, but customizing models to specific needs, addressing bias in specific use cases, and dealing with ethical issues and copyright issues require multi-disciplinary evidence-based research and evaluation. Rudolph et al. [13] suggest exercising caution in adopting a regulatory approach primarily focused on uncovering academic misconduct, such as detecting the use of ChatGPT. Instead, they advocate for cultivating trusting relationships with students in a student-centric pedagogy. Zhai [14] showed that ChatGPT could help researchers write coherent, informative, and systematic papers, and suggests adjusting learning goals—students should be able to use AI tools for scientific research, improving creativity and critical thinking, rather than just focusing on general skills.
@@ III Methods
Remarkably, even with the usage restrictions for ChatGPT and GPT-4 in China, their release has still elicited widespread use and fervent discussions among the public on Weibo, making it the trending topic multiple times. This uncommon situation indicates that people in China are immensely interested in this groundbreaking technological advancement.
To initiate our analysis, we selected key topic keywords in two main areas: AI for Education, and Human-AI Interaction, as illustrated in Figure 1 under “Topic Keywords.” To achieve a comprehensive analysis, we implemented temporal segmentation, dividing the timeline into two distinct periods. The first period covers discussions from the release date of ChatGPT-3 (November 30, 2022) to the eve of GPT-4’s release (March 14, 2023) and the second period begins from the release of GPT-4 (March 15, 2023) onwards. This approach not only enabled us to systematically categorize the collected data but also to trace the evolution of public perception in response to the advancements in AI technology. The WeiboSuperSpider [26] was instrumental in this process, allowing us to efficiently filter and collect posts that matched our designated topic keywords over the specified time frames. By combining these keywords with the #ChatGPT hashtag as our search parameters, we were able to specifically target and analyze the discourse surrounding the integration of AI in education and human-AI interaction on Weibo.
@@
Subsequently, we conducted sentiment analysis on each Weibo post, using the multilingual XLM-roBERTa-base model [30] capable of processing Chinese text. The primary purpose of choosing sentiment analysis is that, as a quantitative tool, it enables us to systematically categorize and analyze subjective information within extensive text data. This approach allows us to methodically assess the emotions expressed in Weibo posts. We opted for a three-category sentiment analysis {positive, negative, neutral} instead of a binary classification {positive, negative}, because we found many social media posts do not convey a clear emotional direction, and a binary classification could fail to accurately reflect the true sentiment of the posts. We then performed an ANOVA analysis [31] on the sentiment analysis results to identify any significant changes in sentiment before and after the launch of GPT-4, shedding light on its impact on public opinion. Additionally, we utilized the LDA model [32] to analyze which topics people were concerned about. To effectively analyze attitudes and perspectives, we retained adjectives and adjectival nouns—words that can convey evaluative meaning. Finally, we combined these results with corresponding Weibo content to provide a comprehensive overview of general attitudes and perspectives on the various scenarios.
@@
Keywords | ChatGPT-3 | GPT-4 |
AI |
29501 |
23547 |
Artificial Intelligence |
35050 |
16471 |
Paper |
8151 |
3190 |
Education |
6076 |
2951 |
New Bing |
2111 |
Law |
1658 |
1929 |
Homework |
4987 |
1485 |
Examination |
1921 |
1325 |
Copilot Office |
998 |
Human-Computer Interaction |
957 |
299 |
Cheating |
1569 |
138 |
Moral Ethics |
204 |
129 |
Answering questions |
119 |
80 |
QA Skills |
63 |
25 |
Originality |
147 |
47 |
Human-Computer Collaboration |
73 |
18 |
@@ IV Results and Analysis
We initially counted the posts and deduplicated user statistics from our scraped data, as detailed in Table III. The data volume meets our research needs and indicates that the usage restrictions on ChatGPT in China have not significantly hindered its discussions on social media platforms (other platforms also have related discussions). Additionally, it’s worth mentioning that 5,780 users discussed both ChatGPT-3 and GPT-4, suggesting that approximately 20,000 users might have been attracted by GPT-4 directly without previous releases. This highlights the fact that the release of GPT-4 attracted the attention of a significant number of Chinese users who were either unfamiliar with it or had limited prior exposure, indicating the appeal of this innovative AI tool.
@@
ChatGPT-3 |
30491 |
92672 |
GPT-4 |
24647 |
78102 |
Total |
49358 |
170774 |
Sentiment | Education | Examination and Cheating | Assisting with Homework | Paper and Originality | Artificial Intelligence | Human-Computer Collaboration and Interaction | Law, Morals and Ethics | |||||||
F |
p |
F |
p |
F |
p |
F |
p |
F |
p |
F |
p |
F |
p |
Negative |
65.6 |
6.1e-16 |
235.7 |
5.2e-52 |
20.0 |
7.8e-06 |
13.1 |
2.8e-04 |
92.8 |
6.0e-22 |
2.4 |
1.2e-01 |
2.1 |
1.4e-01 |
Neutral |
49.6 |
2.1e-12 |
124.4 |
1.5e-28 |
145.0 |
4.5e-33 |
22.8 |
1.9e-06 |
106.6 |
8.8e-37 |
5.8 |
1.6e-02 |
12.6 |
4.0e-04 |
Positive |
24.0 |
9.8e-07 |
267.4 |
1.4e-58 |
508.4 | 1.3e-98 |
109.4 |
1.8e-25 |
0.03 | 8.5e-01 |
19.9 |
8.9e-06 |
6.1 |
1.4e-02 |
@@ IV-A ChatGPT and Education
We notice a near doubling of the proportion of positive attitudes following GPT-4’s release, reaching 4.4, while the proportion of negative attitudes decreases slightly to 12.2. In the ANOVA analysis of the education sector and its three sub-sectors (Table IV), results showed high F-statistics and p-values significantly below 0.05. This indicates that there is a statistically significant difference in the distribution of public sentiment on Weibo before and after the release of GPT-4. This may be because the improvement of GPT-4 over its predecessor have enhanced people’s understanding of ChatGPT’s potential benefits and limitations, leading to increased confidence and optimism among users. Those initially skeptical or hesitant about using ChatGPT might have become more receptive to its application in education, reflecting this shift in attitude.
@@
ChatGPT can generate smooth and clear answers, making it a helpful tool for exams. GPT-4 has even scored in the top 10 on a simulated lawyer exam, showcasing its human-like performance [7]. However, this also raises concerns about potential cheating risks. In exam and cheating scenario, the proportion of individuals who view ChatGPT negatively is significantly higher than those with a positive attitude, reaching 41.9 before the release of GPT-4. This indicates that people are extremely worried that ChatGPT may be used for cheating and undermine the fairness of exams. Following the release of GPT-4, this proportion dropped to 23.7, and the proportion of positive attitudes increased from 2.1 to 5.9. This might be attributed to the increased familiarity with ChatGPT, coupled with schools beginning to implement countermeasures like detection tools to curtail cheating, thereby alleviating concerns and positively influencing attitudes.
@@
Negative |
In the new semester, teachers have started to get a headache about how to prevent students from cheating with ChatGPT writing assignments and reports, which is too difficult. |
Neutral |
I just want to know if I can use ChatGPT to answer during the exam. |
Positive |
#ChatGPT# used 1 month to help my child’s English score improve by a large margin and the test was very easy. |
@@
ChatGPT can play the role of a virtual teacher, helping students understand assignment requirements, improving writing skills, and even providing direct answers. Regarding using ChatGPT for homework assistance, it is worth noting that there has been little change in the proportion of negative attitudes following the release of GPT-4. However, the proportion of positive attitudes increased from 4.7 to 17.9, a growth rate that exceeds all other scenarios. The highest F-statistic and the lowest p-value in the ANOVA analysis results also corroborate this significant change. We believe that this might be due to the fact that using ChatGPT to assist with homework is the most direct and common scenario in which ChatGPT is applied in education. The functionality enhancement of GPT-4, combined with the experience accumulated by users in previous use, has enabled students to have a better experience when using ChatGPT to assist with their homework, thus increasing their confidence and positive evaluations of it. In addition, positive word-of-mouth and recommendations from peers may have a positive impact on students’ attitudes.
@@
Negative |
This software really achieves the ultimate in AI. It is too intelligent, and long-term use will naturally reduce people’s thinking and creativity. If the information provided by ChatGPT is wrong, it will mislead the direction of public opinion. |
Neutral |
ChatGPT is good for homework, but not so feasible for work. |
Positive |
I love chatting with ChatGPT so much, not only can it infer the typos I made, but I can also learn its logic when answering questions and its skills when answering vague questions. |
@@
We can see from the sentiment analysis that although the proportion of negative attitudes is still higher than that of positive attitudes, the relative gap between the two is lower than in the previous two scenarios. The proportion of positive attitudes increased to 15.1 after the release of GPT-4, and it is the only one that exceeds half of the proportion of negative attitudes in our examined education-related scenarios.
@@
Negative |
#chatgpt# This AI is a bit overblown. Its thinking ability is insufficient, and the content it generates lacks originality. It spits out whatever humans feed it. |
Neutral |
I wrote my paper for 3 hours, two and a half of which were spent talking to ChatGPT. |
Positive |
#ChatGPT# I feel that it is very useful for summarizing and refining the main points. It can also give you a general outline to provide ideas for your paper. But a lot of references are made up. I heard that gpt4 had changed this point and is more intelligent. |

### v4

# A Perspective Study on Chinese Social Media regarding LLM for Education and Beyond

@@ Abstract
The application of AI-powered tools has piqued the interest of many fields, particularly in the academic community. This study uses ChatGPT, currently the most powerful and popular AI tool, as a representative example to analyze how the Chinese public perceives the potential of large language models (LLMs) for educational and general purposes. Although facing accessibility challenges, we found that the number of discussions on ChatGPT per month is 16 times that of Ernie Bot developed by Baidu, the most popular alternative product to ChatGPT in the mainland, making ChatGPT a more suitable subject for our analysis. The study also serves as the first effort to investigate the changes in public opinion as AI technologies become more advanced and intelligent. The analysis reveals that, upon first encounters with advanced AI that was not yet highly capable, some social media users believed that AI advancements would benefit education and society, while others feared that advanced AI, like ChatGPT, would make humans feel inferior and lead to problems such as cheating and a decline in moral principles. The majority of users remained neutral. Interestingly, with the rapid development and improvement of AI capabilities, public attitudes have tended to shift in a positive direction. We present a thorough analysis of the trending shift and a roadmap to ensure the ethical application of ChatGPT-like models in education and beyond.
@@
Artificial Intelligence (AI), ChatGPT, Large Language Model (LLM), Chatbots, Education, Ethics, Human-Computer Interaction (HCI).I Introduction
Chatbots can understand natural language input and respond in a human-like manner, making them ideal for tasks such as answering questions, providing guidance, and even offering emotional support. In customer service, chatbots can handle common problems and provide support and guidance. For example, Cui et al. [1] introduced SuperAgent, which is cost-effective when answering repetitive questions, freeing up human support staff to answer more complicated questions. Chatbot can also help users find interesting products and recommend related products in online shopping services. As confirmed by [2], anthropomorphic bot plays a positive role in shaping consumers’ intentions to purchase through chatbot commerce. In the medical field, chatbots serve as medical manuals to help patients become aware of their illnesses and improve their health. For example, a text diagnosis bot enables sufferers to join in analyses of their medicinal matters and present a personalized analysis report about the symptoms [3]. Meanwhile, research on the application of chatbots to education and other areas is still in its infancy [4] and waits for more exploration.
@@
To address these research questions, we collected data by crawling posts containing “ChatGPT” and other relevant keywords on Sina Weibo, China’s largest microblogging platform with 605 million monthly active users. We chose ChatGPT over more easily accessible AI products within China for several reasons.
Firstly, we emphasize that even though there are usage restrictions for ChatGPT in China, alternative access methods are available. Additionally, numerous online videos demonstrate the capabilities and functionalities of GPT, which help users understand and appreciate its potential.
Secondly, Our research primarily focuses on the general awareness and understanding of AI-powered tools among the public, not just those with hands-on experience. Understandings can be acquired through various sources such as media reports, academic discussions, and social networks, which are equally important and effective. This approach is common in social science research, particularly in the early stages of technological adoption such as Self-driving automobiles.
As a comparison, we also crawled discussion posts about Ernie Bot, an AI chatbot developed by Baidu, which is the most popular and representative ChatGPT alternative in China. As shown in Table IV, there were only 25,462 posts about Ernie Bot from its release on 16 March 2023 to 17 May 2024, compared to 145,219 posts about ChatGPT in the five months from 30 November 2022 to 4 May 2023. The discussions about ChatGPT per month were approximately 16 times more frequent than those about Ernie Bot, emphasizing the significant social media presence and impact of ChatGPT. Therefore, analyzing ChatGPT, a remarkable AI application, is more appropriate for our study.
| Papers | Brief Introduction |
| Mogavi et al. [10] | This article analyzed data from , Reddit, YouTube, and LinkedIn to explore the user experience and perspectives of early adopters toward ChatGPT in various education sectors. |
| Baidoo-Anu et al. [11] | This review synthesized recent extant literature to offer some potential benefits of ChatGPT in promoting teaching and learning. |
| Kasneci et al. [12] | This commentary discusses the potential benefits and challenges of LLMs in education from student and teacher perspectives. |
| Rudolph et al. [13] | This review adopts a desktop analysis approach and conducts an extensive literature review, focusing on ChatGPT’s implications for higher education. |
| Zhai [14] | This study piloted ChatGPT to write an academic paper and reflected on the potential impacts of AI tools on education based on the user experience. |
| Tlili et al. [15] | This study collected and analyzed data, and examined ChatGPT in education among early adopters through a qualitative instrumental case study. |
| Kung et al. [16] | This article evaluated the performance of ChatGPT on the United States Medical Licensing Exam and suggested that LLMs may have the potential to assist with medical education and clinical decision-making. |
| Rahman et al. [17] | This study analyzed data from articles, websites, blogs, and various media, finding ChatGPT effective in generating initial academic research ideas. However, it identified challenges in literature synthesis, citations, problem statements, research gaps, and data analysis, suggesting the need for guidelines on the proper use of LLMs. |
| Bahrini et al. [18] | This article reviewed the existing literature, and examined the applications, opportunities, and threats of ChatGPT in 10 main domains, providing detailed examples for the business and industry as well as education. |
| Liu et al. [19] | This article performed an in-depth analysis of 194 relevant papers on arXiv, and the findings reveal a significant and increasing interest in ChatGPT/GPT-4 research. |
| Our article | Our article differs in the following ways: 1) we collect data from Chinese social media, Weibo, which serves as a unique public opinion environment and data source of Chinese society that has not been explored; 2) we compare the people’s attitudes and opinions before and after the release of GPT-4; 3) we expand the investigation to more interactions scenarios beyond education. |
After getting the data, we categorized the posts gathered by the specified keywords into two major groups: AI for education and human-AI interaction. Each major category includes three scenarios. We conducted sentiment and topic analyses on these posts to assess the positive and negative impacts of ChatGPT, aiming to enhance its benefits while mitigating potential drawbacks for its safe and responsible use.
@@
Public concerns include privacy infringement, content bias, inaccuracy, harm to educational fairness, and over-reliance on AI applications.
@@ II Related Work
Education is one of the most widely engaging application fields for ChatGPT, with a broad audience and easy deployment. Mogavi et al. [10] analyzed social media data from , Reddit, YouTube, and LinkedIn, revealing that the public discourse is generally positive, and there is enthusiasm regarding ChatGPT’s use in education. Nevertheless, they found that an overreliance on AI systems may encourage superficial learning habits and diminish students’ social and critical thinking skills. Baidoo-Anu et al. [11] highlight ChatGPT’s capability to assess students’ learning styles, offer personalized tutoring and feedback, and even grade assignments, thereby freeing up teachers for other tasks. However, concerns about the absence of human interaction in such generative models are noted, especially for students needing personal teacher engagement for an optimal learning experience. Kasneci et al. [12] believe that ChatGPT can provide customized services for learners at different stages and fields, but customizing models to specific needs, addressing bias in specific use cases, and dealing with ethical issues and copyright issues require multi-disciplinary evidence-based research and evaluation. Rudolph et al. [13] suggest exercising caution in adopting a regulatory approach primarily focused on uncovering academic misconduct, such as detecting the use of ChatGPT. Instead, they advocate for cultivating trusting relationships with students in a student-centric pedagogy. Zhai [14] showed that ChatGPT could help researchers write coherent, informative, and systematic papers, and suggests adjusting learning goals—students should be able to use AI tools for scientific research, improving creativity and critical thinking, rather than just focusing on general skills.
@@ III Methods
To initiate our analysis, we selected key topic keywords in two main areas: AI for Education, and Human-AI Interaction, as illustrated in Figure 1 under “Topic Keywords.” To achieve a comprehensive analysis, we implemented temporal segmentation, dividing the timeline into two distinct periods. The first period covers discussions from the release date of ChatGPT-3 (November 30, 2022) to the eve of GPT-4’s release (March 14, 2023) and the second period begins from the release of GPT-4 (March 15, 2023) to May 4, 2023. This approach not only enabled us to systematically categorize the collected data but also to trace the evolution of public perception in response to the advancements in AI technology. The WeiboSuperSpider [26] was instrumental in this process, allowing us to efficiently filter and collect posts that matched our designated topic keywords over the specified time frames. By combining these keywords with the #ChatGPT hashtag as our search parameters, we were able to specifically target and analyze the discourse surrounding the integration of AI in education and human-AI interaction on Weibo.
@@
| Keywords | ChatGPT | ChatGPT | Ernie Bot |
| 22/11/30-23/03/14 | 23/03/15-23/05/04 | 23/03/16-24/05/17 | |
| AI | 29501 | 23547 | 15617 |
| Artificial Intelligence | 35050 | 16471 | 6287 |
| Paper | 8151 | 3190 | 730 |
| Education | 6076 | 2951 | 1091 |
| New Bing | - | 2111 | - |
| Law | 1658 | 1929 | 569 |
| Homework | 4987 | 1485 | 390 |
| Examination | 1921 | 1325 | 470 |
| Copilot Office | - | 998 | - |
| Human-Computer Interaction | 957 | 299 | 140 |
| Cheating | 1569 | 138 | 39 |
| Moral Ethics | 204 | 129 | 4 |
| Answering questions | 119 | 80 | 96 |
| QA Skills | 63 | 25 | 0 |
| Originality | 147 | 47 | 13 |
| Human-Computer Collaboration | 73 | 18 | 16 |
@@
Subsequently, we conducted sentiment analysis on each Weibo post, using the multilingual XLM-roBERTa-base model [30] capable of processing Chinese text. The primary purpose of choosing sentiment analysis is that, as a quantitative tool, it enables us to systematically categorize and analyze subjective information within extensive text data. This approach allows us to methodically assess the emotions expressed in Weibo posts. We opted for a three-category sentiment analysis {positive, negative, neutral} instead of a binary classification {positive, negative}, because we found many social media posts do not convey a clear emotional direction, and a binary classification could fail to accurately reflect the true sentiment of the posts. We then performed a chi-square test [31] on the sentiment analysis results to identify any significant changes in sentiment before and after the launch of GPT-4, shedding light on its impact on public opinion. Additionally, we utilized the LDA model [32] to analyze which topics people were concerned about. To effectively analyze attitudes and perspectives, we retained adjectives and adjectival nouns—words that can convey evaluative meaning. Finally, we combined these results with corresponding Weibo content to provide a comprehensive overview of general attitudes and perspectives on the various scenarios.
@@
| Sentiment | Education | Examination and Cheating | Assisting with Homework | Paper and Originality | ||||
| p | p | p | p | |||||
| Positive | ||||||||
| Neutral | ||||||||
| Negative | ||||||||
@@ IV Results and Analysis
We initially counted the posts and deduplicated user statistics from our scraped data, as detailed in Table IV. The data volume meets our research needs and indicates that the usage restrictions on ChatGPT in China have not significantly hindered its discussions on social media platforms (other platforms also have related discussions). Additionally, it’s worth mentioning that 5,780 users discussed both ChatGPT-3 and GPT-4, suggesting that approximately 20,000 users might have been attracted by GPT-4 directly without previous releases. This highlights the fact that the release of GPT-4 attracted the attention of a significant number of Chinese users who were either unfamiliar with it or had limited prior exposure, indicating the appeal of this innovative AI tool.
@@
| ChatGPT Period 1 | 30491 | 90476 |
| ChatGPT Period 2 | 24647 | 54743 |
| ChatGPT Total | 49358 | 145219 |
| Ernie Bot | 10619 | 25462 |
@@ IV-A ChatGPT and Education
We notice a near doubling of the proportion of positive attitudes following GPT-4’s release, reaching 4.4, while the proportion of negative attitudes decreases slightly to 12.2. In the chi-square test of the education sector and its three sub-sectors (Table III), most results showed high chi-square statistics and p-values significantly below 0.05. This indicates that there is a statistically significant difference in the distribution of public sentiment on Weibo before and after the release of GPT-4. This may be because the improvement of GPT-4 over its predecessor have enhanced people’s understanding of ChatGPT’s potential benefits and limitations, leading to increased confidence and optimism among users. Those initially skeptical or hesitant about using ChatGPT might have become more receptive to its application in education, reflecting this shift in attitude.
@@
ChatGPT can generate smooth and clear answers, making it a helpful tool for exams. GPT-4 has even scored in the top 10 on a simulated lawyer exam, showcasing its human-like performance [7]. However, this also raises concerns about potential cheating risks. In exam and cheating scenario, the proportion of individuals who view ChatGPT negatively is significantly higher than those with a positive attitude, reaching 41.9 before the release of GPT-4. This indicates that people are extremely worried that ChatGPT may be used for cheating and undermine the fairness of exams. Following the release of GPT-4, this proportion dropped to 23.7, and the proportion of positive attitudes increased from 2.1 to 5.9, Following the release of GPT-4, this proportion dropped to 23.7%, and the proportion of positive attitudes increased from 2.1% to 5.9%, which is confirmed by the chi-square test results, with for positive sentiment and for negative sentiment. This might be attributed to the increased familiarity with ChatGPT, coupled with schools beginning to implement countermeasures like detection tools to curtail cheating, thereby alleviating concerns and positively influencing attitudes.
@@
| Negative | In the new semester, teachers have started to get a headache about how to prevent students from cheating with ChatGPT writing assignments and reports, which is too difficult. |
| Neutral | I just want to know if I can use ChatGPT to answer during the exam. |
| Positive | #ChatGPT# used 1 month to help my child’s English score improve by a large margin and the test was very easy. |
@@
ChatGPT can play the role of a virtual teacher, helping students understand assignment requirements, improving writing skills, and even providing direct answers. Regarding using ChatGPT for homework assistance, it is worth noting that there has been little change in the proportion of negative attitudes following the release of GPT-4. However, the proportion of positive attitudes increased from 4.7 to 17.9, a growth rate that exceeds all other scenarios. The high chi-squared statistics and low p-values across the three sentiments also corroborate this significant change. We believe that this might be due to the fact that using ChatGPT to assist with homework is the most direct and common scenario in which ChatGPT is applied in education. The functionality enhancement of GPT-4, combined with the experience accumulated by users in previous use, has enabled students to have a better experience when using ChatGPT to assist with their homework, thus increasing their confidence and positive evaluations of it. In addition, positive word-of-mouth and recommendations from peers may have a positive impact on students’ attitudes.
@@
| Negative | This software really achieves the ultimate in AI. It is too intelligent, and long-term use will naturally reduce people’s thinking and creativity. If the information provided by ChatGPT is wrong, it will mislead the direction of public opinion. |
| Neutral | ChatGPT is good for homework, but not so feasible for work. |
| Positive | I love chatting with ChatGPT so much, not only can it infer the typos I made, but I can also learn its logic when answering questions and its skills when answering vague questions. |
@@
We can see from the sentiment analysis that although the proportion of negative attitudes is still higher than that of positive attitudes, the relative gap between the two is lower than in the previous two scenarios. The proportion of positive attitudes increased to 15.1 after the release of GPT-4 with and , and it is the only one that exceeds half of the proportion of negative attitudes in our examined education-related scenarios.
@@
| Negative | #chatgpt# This AI is a bit overblown. Its thinking ability is insufficient, and the content it generates lacks originality. It spits out whatever humans feed it. |
| Neutral | I wrote my paper for 3 hours, two and a half of which were spent talking to ChatGPT. |
| Positive | #ChatGPT# I feel that it is very useful for summarizing and refining the main points. It can also give you a general outline to provide ideas for your paper. But a lot of references are made up. I heard that gpt4 had changed this point and is more intelligent. |